### v1

# Simulating discrete-time quantum walk with urn model

Abstract
Urn models have long been used to study stochastic processes, probability distributions, and reinforcement dynamics. Meanwhile, discrete time quantum walks (DTQW) serve as fundamental components in quantum computation and quantum information theory. This work explores a novel connection between an urn model and discrete-time quantum walks, focusing on how urn-based stochastic processes can provide insights into quantum state evolution and algorithmic behavior. Importantly, our model operates entirely with real-valued quantities, without relying on complex amplitudes of the quantum walk. Every stochastic rule is derived from urn-based processes, whose parameters are governed by the coin and shift mechanisms of the quantum walk, maintaining the structural dynamics of the QW. Through this interdisciplinary approach, we aim to bridge classical and quantum frameworks, offering new perspectives on both quantum algorithms and stochastic modeling. This work connects the urn model to the broad range of applications where DTQWs are successfully employed.
Exploring the dynamics of a physical system often involves identifying suitably engineered dynamics in a potentially very different realm that exhibits analogous behavior. This approach not only enhances our understanding of the physical system but also enriches the dynamics of that different realm, making both perspectives more insightful and beneficial. A classic example emerged in 2002, when Grover’s quantum search algorithm was shown to correspond to the behavior of coupled pendulums [1]. Recent work [2] clarified this connection further. It showed that the dynamics of any classical system of harmonic oscillators can be equivalently understood as the dynamics of a corresponding quantum system of exponentially smaller size. Conversely, every quantum algorithm can be equivalently understood as the propagation of a classical wave in a system of coupled oscillators, albeit one that may involve action at a distance and significant classical resources, which make the latter process impractical. But the important point is that such an analogy led to the discovery of an efficient quantum algorithm to simulate the classical dynamics of coupled oscillators.
This duality underscores the importance of systematically studying how useful classical systems—such as coupled oscillators, which can themselves simulate a wide variety of classical dynamics—can reflect the behavior of quantum algorithms. While such classical simulations might incur substantial computational costs or non-local characteristics, they can lead to the development of novel quantum algorithms for simulating intricate classical systems, as well as quantum-inspired classical strategies. In this spirit, we employ discrete-time quantum walks (DTQW) and construct an urn model as the corresponding useful classical system, offering a fresh route to replicate quantum dynamics.
Quantum walks on graphs are one of the few established design strategies for quantum algorithms. Despite their simple description, quantum walks are extremely powerful, as they are universal for quantum computation [3, 4]. Owing to their non-classical spreading dynamics, localization property, and potential algorithmic advantages, quantum walks have been extensively explored both theoretically and experimentally [5, 6, 7].
On the other hand, Urn models—built on the simple act of drawing and replacing colored balls—serve as a versatile framework for modeling both random and self‑reinforcing dynamics across domains as varied as population genetics, network growth, statistical physics, and machine learning [8, 9].
Replicating quantum phenomena like DTQWs on useful classical platforms is inherently challenging. This difficulty stems from the fact that DTQWs rely on quantum superposition, where evolving states interfere—constructively or destructively—based on their relative phases. As a result, they produce probability distributions that differ from those of classical random walks.
A line of research on this topic has approached this challenge by constructing classical random walks (CRWs) [10, 11] whose space and time-dependent transition probabilities in each step are extracted from the complex amplitudes of the DTQW. This approach yields a non-homogeneous CRW whose probability distribution replicates that of the corresponding DTQW. Although effective in reproducing certain statistics, these methods necessarily break the spatial and temporal structure of the DTQW and demand access to quantum amplitudes at each step, an approach analogous to Bohmian mechanics[12], where the quantum wave function guides particle trajectories.
In contrast, alternative formulations of quantum mechanics—akin to the hydrodynamic approaches of Holland and Poirier [13, 14, 15, 16, 17, 18] —eschew direct reliance on the wave function, instead constructing real-valued fields or potentials that reproduce quantum evolution without amplitude-based guidance. Inspired by this distinction, we propose an urn‑model framework for simulating DTQWs that never invokes complex amplitudes. All transition rules arise from real‑valued urn processes whose parameters depend only on the QW’s coin and shift structure. Crucially, our model mirrors the structural properties of the original quantum walk system. When the QW has uniform coin and shift operations, our urn model remains homogeneous in space and time, employing identical urn parameters at each node and time step. If QW coin or shift operations vary across space or time, the corresponding urn model adapts its parameters in the same pattern, preserving the structural variations of QW. This yields two main advantages. First, preserving the structural variations of the QW simplifies both the analysis and the simulation, while allowing our method to be naturally extended to higher-dimensional graphs. Second, by avoiding amplitude-based guidance, the urn model brings into focus a fundamental probabilistic mechanism underlying quantum walk spreading and interference, insights that may carry foundational significance.
To define the one-dimensional discrete-time quantum walk we require the coin Hilbert space $\mathcal{H}_{c}$ and the position Hilbert space $\mathcal{H}_{p}$. The $\mathcal{H}_{c}$ is spanned by the internal basis states of the particle, $|0\rangle$ and $|1\rangle$, and the $\mathcal{H}_{p}$ is spanned by the position basis states $|n\rangle$, $n\in\mathbb{Z}$. The total system is then in the space $\mathcal{H}=\mathcal{H}_{c}\otimes\mathcal{H}_{p}$. To implement a simple version of the quantum walk, known as the Hadamard walk, one applies the Hadamard operator
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\[4.0pt] 1&-1\end{pmatrix}$ (1)
on $\mathcal{H}_{c}$, which transforms a particle at the origin in one of the basis states into an equal superposition of $\lvert 0\rangle$ and $\lvert 1\rangle$. Next, we perform the conditional shift
$S=\bigl{(}\lvert 0\rangle\langle 0\rvert\otimes\sum_{n=-\infty}^{\infty}\lvert n+1\rangle\langle n\rvert\bigr{)}+\bigl{(}\lvert 1\rangle\langle 1\rvert\otimes\sum_{n=-\infty}^{\infty}\lvert n-1\rangle\langle n\rvert\bigr{)}.$ (2)
A single step of the Hadamard quantum walk is thus
$W\;=\;S\bigl{(}H\otimes\mathbb{I}_{p}\bigr{)},$ (3)
The process of $W$ is iterated without resorting to the intermediate measurements to realize a large number of steps of the quantum walk.
It has been demonstrated in [19] that a three-parameter $SU(2)$ quantum coin operation,
$C(\xi,\theta,\zeta)=\begin{pmatrix}e^{i\xi}\cos\theta&e^{i\zeta}\sin\theta\\[6.0pt] -e^{-i\zeta}\sin\theta&e^{-i\xi}\cos\theta\end{pmatrix},$ (4)
is sufficient to describe the most general form of the discrete-time quantum walk, where $\xi,\theta,\zeta\in\mathbb{R}$ and $\theta\in[0,\pi/2]$. By choosing different $(\xi,\theta,\zeta)$, one obtains a variety of discrete-time walks with distinct dynamical features.
Consider a general single-qubit quantum state given by
$\lvert\psi\rangle=\psi_{0}\lvert 0\rangle+\psi_{1}\lvert 1\rangle.$ (5)
Here, the complex amplitudes can be written as $\psi_{0}=r_{0}e^{i\phi_{0}}$ and $\psi_{1}=r_{1}e^{i\phi_{1}}$. The probabilities of measuring the qubit in the computational basis states $\lvert 0\rangle$ and $\lvert 1\rangle$ are given by
$\begin{array}[]{rc}P_{0}=\mid\psi_{0}\mid^{2}=r_{0}^{2}\\ P_{1}=\mid\psi_{1}\mid^{2}=r_{1}^{2}\end{array}$ (8)
Applying the Hadamard gate to the state $\lvert\psi\rangle$ transforms it as follows:
$\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\begin{bmatrix}\psi_{0}\\ \psi_{1}\end{bmatrix}=\frac{1}{\sqrt{2}}\begin{bmatrix}\psi_{0}+\psi_{1}\\ \psi_{0}-\psi_{1}\end{bmatrix}=\begin{bmatrix}\psi_{0}^{\prime}\\ \psi_{1}^{\prime}\end{bmatrix}$ (9)
Suppose $\psi_{0}^{\prime}=r_{0}^{\prime}e^{i\phi_{0}^{\prime}}$ and $\psi_{1}^{\prime}=r_{1}^{\prime}e^{i\phi_{1}^{\prime}}$ then one can write
$\begin{array}[]{rc}P_{0}^{\prime}={r_{0}^{\prime}}^{2}=\frac{1}{2}(r_{0}^{2}+r_{1}^{2}+2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})),\\ P_{1}^{\prime}={r_{1}^{\prime}}^{2}=\frac{1}{2}(r_{0}^{2}+r_{1}^{2}-2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})),\end{array}$ (12)
and
$\begin{array}[]{rc}\phi_{0}^{\prime}=\tan^{-1}\left(\frac{r_{0}\sin\phi_{0}+r_{1}\sin\phi_{1}}{r_{0}\cos\phi_{0}+r_{1}\cos\phi_{1}}\right),\\ \phi_{1}^{\prime}=\tan^{-1}\left(\frac{r_{0}\sin\phi_{0}-r_{1}\sin\phi_{1}}{r_{0}\cos\phi_{0}-r_{1}\cos\phi_{1}}\right).\end{array}$ (15)
Where $P_{0}^{\prime}$ and $P_{1}^{\prime}$ are resulting probabilities of measuring the transformed state in the computational basis. We now construct an urn model specifically designed to replicate the measurement statistics of a single-qubit state (Eq. 8), as well as the statistics following the application of a Hadamard operation on that state (Eq. 12 and 15).
An urn is a container that holds a collection of objects, usually called balls. The contents of the urn can be altered by randomly removing or adding balls according to the rules given. The behavior of such a system is probabilistic in nature. In the literature, it has been noted earlier that urn models has the capability to describe various quantum logic [20, 21].
Imagine a setting with two urns, labeled 00 and $1$, and a total collection of $N$ balls. The various operations on this system are categorized into three classes: preparation ($\mathbb{P}$), transformation ($\mathbb{T}$), and measurement ($\mathbb{M}$).
In a typical $\mathbb{P}$ procedure, we begin by selecting one ball uniformly at random from the collection of $N$ balls and marking it for future use. Following this, the $N$ balls (including the marked ball) are distributed between the two urns. The allocation can proceed in one of two ways, each chosen with equal probability:
  • Either $N_{0}=\lfloor\rho_{0}N\rfloor$ balls are randomly selected from the collection and placed into urn 00, with the remaining $N_{1}=N-N_{0}$ balls allocated to urn 1;
  • Or $N_{0}=\lceil\rho_{0}N\rceil$ balls are randomly selected from the collection and placed into urn 0, with the remaining $N_{1}=N-N_{0}$ balls placed into urn 1.
Here, the parameter $\rho_{0}\in[0,1]$ governs the expected proportion of balls assigned to urn 00. It is important to note that, during the selection procedure, all balls—including the marked one—are treated as identical. Additionally, each urn $i$ (for $i=0,1$) is assigned a real-valued phase tag $\eta_{i}$. The prepared state is denoted by $\textit{U}^{N,N_{0}}_{\eta_{0},\eta_{1}}$.
Now, let us define the transformation process $\mathbb{T}$ on the state $\textit{U}^{N,N_{0}}_{\eta_{0},\eta_{1}}$. In the $\mathbb{T}$ process, some randomly selected balls from one of the urns are transferred to the other urn according to a certain stochastic rule. As a result, the number of balls in each urn is updated and becomes either $N_{0}^{\prime}=\lfloor\tilde{N}_{0}\rfloor$ and $N_{1}^{\prime}=N-\lfloor\tilde{N}_{0}\rfloor$, or $N_{0}^{\prime}=\lceil\tilde{N}_{0}\rceil$ and $N_{1}^{\prime}=N-\lceil\tilde{N}_{0}\rceil$, each with equal probability. Here, $\tilde{N}_{0}$ is a function of $N_{0},N_{1},\eta_{0}$ and $\eta_{1}$.
The phase tags $\eta_{0}$ and $\eta_{1}$ are also updated to new values $\eta_{0}^{\prime}$ and $\eta_{1}^{\prime}$, where $\eta_{i}^{\prime}=\eta_{i}^{\prime}(N_{0},N_{1},\eta_{0},\eta_{1})$.
Putting it all together, the transformation is described as:
$\mathbb{T}:\textit{U}^{N,N_{0}}_{\eta_{0},\eta_{1}}\longrightarrow\textit{U}^{N,N_{0}^{\prime}}_{\eta_{0}^{\prime},\eta_{1}^{\prime}}$
Now, let us define a process referred to as the measurement process, denoted by $\mathbb{M}$. In the $\mathbb{M}$ process, all balls are placed into the urn that contains the marked ball, and the measurement outcome is the number associated with that urn. Both phase tags are then reset to zero.
Thus, the post-measurement state becomes:
  • $U^{N,N}_{0,0}$ if the marked ball is in urn 0;
  • $U^{N,0}_{0,0}$ if the marked ball is in urn 1.
For example, suppose that after a $\mathbb{T}$ process the marked ball belongs to urn 1. Then, the measurement outcome will be 1, and all $N$ balls will be placed into urn 1. If the $\mathbb{M}$ process is applied again to the urn system, the outcome will remain the same, because the marked ball will always be found in urn 1.
Now, suppose that $\mathbb{P}$ and $\mathbb{M}$ are applied to the system without any $\mathbb{T}$ process in between. In the $\mathbb{P}$ process, we set $\eta_{i}=\phi_{i}$ and $\rho_{0}=r_{0}^{2}$. The probability of randomly selecting a particular ball to be marked from $N$ identical balls is $\frac{1}{N}$. As a result, the probability of obtaining outcomes 0 and 1 after the $\mathbb{M}$ process, for a particular configuration, will be $\frac{N_{0}}{N}$ and $\frac{N_{1}}{N}$, respectively.
Thus, the ensemble average of the probabilities corresponding to outcomes 0 and 1 can be written as:
$P_{0}=\left\langle\frac{N_{0}}{N}\right\rangle=\frac{\lfloor r_{0}^{2}N\rfloor+\lceil r_{0}^{2}N\rceil}{2N}$
and
$P_{1}=1-\left\langle\frac{N_{0}}{N}\right\rangle$
respectively.
Now, one can always write:
$r_{0}^{2}N=\lfloor r_{0}^{2}N\rfloor+\epsilon\quad\text{and}\quad r_{0}^{2}N=\lceil r_{0}^{2}N\rceil-\epsilon^{\prime}$
where $10\leq\epsilon,\epsilon^{\prime}<$ 1. Moreover, whenever $r_{0}^{2}N\notin\mathbb{Z}$, it follows that $\epsilon+\epsilon^{\prime}=1$.
Hence,
$\lfloor r_{0}^{2}N\rfloor+\lceil r_{0}^{2}N\rceil=2r_{0}^{2}N+(\epsilon^{\prime}-\epsilon)$
and thus,
$\lim_{N\to\infty}\frac{\lfloor r_{0}^{2}N\rfloor+\lceil r_{0}^{2}N\rceil}{2N}=\lim_{N\to\infty}\left(r_{0}^{2}+\frac{\epsilon^{\prime}-\epsilon}{2N}\right)=r_{0}^{2}$
Therefore, in the large $N$ limit, we obtain $P_{0}=r_{0}^{2}$ and $P_{1}=1-r_{0}^{2}=r_{1}^{2}$.
Now, let us define a specific transformation denoted by $\mathbb{T}_{H}$ acting on $\textit{U}^{N,N_{0}}_{\phi_{0},\phi_{1}}$. This transformation is constructed such that when followed by a $\mathbb{M}$ operation, it produces measurement statistics that converge to the Hadamard operation (Eq. (12)) in the limit $N\to\infty$. Furthermore, the updated phase tags also converge to $\phi_{0}^{\prime}$ and $\phi_{1}^{\prime}$ (as given by Eq. (15)) in that limit.
So, the new quantity $\tilde{N}_{0}$ in the transformation $\mathbb{T}_{H}$ must be chosen such that, as $N\to\infty$, the probability of obtaining the measurement outcome 00 becomes
$\lim_{N\to\infty}\frac{\langle N_{0}^{\prime}\rangle}{N}=\lim_{N\to\infty}\frac{\lfloor\tilde{N}_{0}\rfloor+\lceil\tilde{N}_{0}\rceil}{2N}={r_{0}^{\prime}}^{2},$ (16)
and the probability of outcome $1$ becomes
$\lim_{N\to\infty}\left(1-\frac{\langle N_{0}^{\prime}\rangle}{N}\right)={r_{1}^{\prime}}^{2}.$ (17)
Furthermore, we require
$\lim_{N\to\infty}\langle\eta_{i}^{\prime}(N_{0},N_{1},\phi_{0},\phi_{1})\rangle={\phi_{i}}^{\prime}$ (18)
where $\eta_{i}^{\prime}$ represents the updated phase tags.
To implement this, we adopt the following transformation scheme [see appendix]:
$\tilde{N}_{0}=\frac{N_{0}+N_{1}}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0}).$ (19)
The updated phase tags are given by
$\begin{array}[]{rc}\eta_{i}^{\prime}(N_{0},N_{1},\phi_{0},\phi_{1})=\tan^{-1}\left(\frac{\sqrt{\frac{N_{0}}{N}}\sin\phi_{0}+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\sin\phi_{1}}{\sqrt{\frac{N_{0}}{N}}\cos\phi_{0}+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\cos\phi_{1}}\right)\end{array}$ (21)
According to this scheme, the updated number of balls in urn 0, denoted by $N_{0}^{\prime}$, will either be $\lfloor\frac{N}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0})\rfloor$ or $\lceil\frac{N}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0})\rceil$, each chosen with equal probability. If $N_{0}^{\prime}>N_{0}$, the $\mathbb{T}_{H}$ adds $N_{0}^{\prime}-N_{0}$ balls to urn 0 from urn 1. Conversely, if $N_{0}^{\prime}<N_{0}$, it moves $N_{0}-N_{0}^{\prime}$ balls from urn 0 to urn 1.
One can show that [see appendix] in the limit $N\to\infty$, the measurement outcome probabilities converge to those of the Hadamard transformation.
For finite $N$, there will always be some error, but by taking $N$ large enough, one can approximate such quantum operations with arbitrary precision.
Moreover, this method can be generalized to simulate other single-qubit quantum coin operations, as expressed in Eq. (4). One can obtain the generalized forms of $\tilde{N}_{0}$ and $\eta_{i}^{\prime}$ in the corresponding transformation process denoted by $\mathbb{T}_{c}$ as follows (see appendix):
$\begin{array}[]{rc}\tilde{N}_{0}=N_{0}\cos^{2}\theta+N_{1}\sin^{2}\theta&\\ &+2\sqrt{N_{0}N_{1}}\sin\theta\cos\theta\cos(\phi_{0}-\phi_{1}+\xi-\zeta)\\ \end{array}$ (24)
$\begin{array}[]{rc}\eta_{i}^{\prime}=\tan^{-1}\left(\frac{\sqrt{\frac{N_{0}}{N}}\cos\theta\sin(\xi+\phi_{0})+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\sin\theta\sin(\zeta+\phi_{1})}{\sqrt{\frac{N_{0}}{N}}\cos\theta\cos(\xi+\phi_{0})+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\sin\theta\cos(\zeta+\phi_{1})}\right)\end{array}$ (26)
Finally, we define the urn model corresponding to a DTQW using the generalized $SU(2)$ coin (Eq. (4)), with the initial state given by:
Figure 1: Probability distributions arising from the $\mathbb{M}$ process in the urn model of the Hadamard coined quantum walk, with initial state $\frac{1}{\sqrt{2}}\lvert 0\rangle-\frac{i}{\sqrt{2}}\lvert 1\rangle$ and at time $t=100$, are shown for different total numbers of balls $N$. As $N$ increases, the distribution converges to that of the ideal Hadamard quantum walk.
$\lvert\psi(x,t=0)\rangle=\left(\psi_{0}\lvert 0\rangle+\psi_{1}\lvert 1\rangle\right)\otimes\lvert x=0\rangle$ (27)
We perform the $\mathbb{P}$, $\mathbb{T}_{c}$, and $\mathbb{M}$ processes within a generalized framework. We begin with a collection of $N$ balls. To each lattice site $x\in\mathbb{Z}$, we assign two urns, labeled $(x,0)$ and $(x,1)$. One of the $N$ balls is randomly selected and marked, after which all $N$ balls are distributed among the urns.
The state of the system at time $t$ is fully specified by the nonnegative integers
$N_{x,c}(t)\;=\;\text{number of balls in urn }(x,c),\quad c\in\{0,1\},$
together with associated “phase” tags $\eta_{x,c}(t)\in[0,2\pi)$ for each occupied urn. These obey the global constraint
$\sum_{x\in\mathbb{Z}}\sum_{c=0}^{1}N_{x,c}(t)\;=\;N,\qquad\forall\,t.$ (28)
At time $t=0$, only the two urns at the origin carry balls; all others are empty:
$N_{x,c}(0)=0\quad\text{for}\quad(x,c)\neq(0,0),(0,1).$
Within the origin urns, we set
$N_{0,0}(0)=\begin{cases}\lfloor r_{0}^{2}N\rfloor,&\text{with probability }1/2,\\[6.0pt] \lceil r_{0}^{2}N\rceil,&\text{with probability }1/2,\end{cases}$
$N_{0,1}(0)=N-N_{0,0}(0),$
and assign phases $\eta_{0,0}(0)=\phi_{0}$, $\eta_{0,1}(0)=\phi_{1}$ so as to encode the initial coin–position state of Eq.27.
Now for each discrete time step $t\to t+1$, we perform two sequential sub‐steps:
Figure 2: Snapshot of the trajectories of the marked ball in the urn model ($N=10^{9}$) of the 1D Hadamard quantum walk, initialized in the state $\frac{1}{\sqrt{2}}\lvert 0\rangle-\frac{i}{\sqrt{2}}\lvert 1\rangle$, shown at time $t=200$
Coin–Transformation $\mathbb{T}_{c}$:
At every site $x$ with $N_{x,0}(t)+N_{x,1}(t)>0$, we apply a $\mathbb{T}_{c}$ process on the two urns labeled by $(x,0)$ and $(x,1)$ with $\tilde{N}$ and $\eta_{i}^{\prime}$ as given in Eq. (24) and Eq. (26) respectively. This is a stochastic mixing rule of the two urns at each site that mirrors the quantum “coin” operation. Concretely, urn‐occupancies and phases are updated to new values $N^{\prime}_{x,c}(t)$ and $\eta^{\prime}_{x,c}(t)$ so as to reproduce, in the large–$N$ limit, the SU(2) coin operation of Eq. (4).
Conditional Shift $\mathbb{S}$:
After the $\mathbb{T}_{c}$, all balls in urn $c=0$ at site $x$ move to urn 00 of site $x+1$, while those in $c=1$ move to urn $1$ of site $x-1$. Phases are also carried along unchanged:
$N_{x+1,0}(t+1)$ $=N^{\prime}_{x,0}(t),$ $\eta_{x+1,0}(t+1)$ $=\eta^{\prime}_{x,0}(t),$ (29)
$N_{x-1,1}(t+1)$ $=N^{\prime}_{x,1}(t),$ $\eta_{x-1,1}(t+1)$ $=\eta^{\prime}_{x,1}(t),$ (30)
and all other urn‐occupancies reset to zero.
The $\mathbb{M}$ process is then implemented within this generalized framework as follows.
Measurement ($\mathbb{M}$ process):
After $t$ steps of iterating $\mathbb{T}_{c}$ and $\mathbb{S}$, we perform the measurement process $\mathbb{M}$: we identify the urn $(x_{\rm m},c_{\rm m})$ containing the marked ball, move all balls into that urn, and record the site index $x_{\rm m}$. The resulting histogram of recorded $x_{\rm m}$ over many trials reproduces the probability distribution of the quantum walk at time $t$ in the large $N$ limit [Fig. 1].
The dynamics can be viewed from a different perspective also: by iterating $\mathbb{T}_{c}$ and $\mathbb{S}$ for $t$ steps, the marked ball undergoes a classical random walk [Fig. 2] whose large–$N$ distribution converges to that of the quantum walker’s position distribution $|\psi(x,t)|^{2}$.
In this work, we have developed an urn model to simulate the dynamics of 1D discrete-time quantum walks with a generalized coin. The urn model can be systematically generalized to higher-dimensional graphs. In the two-dimensional case, for example, the walker requires a $4\times 4$ coin operation acting on a two-qubit state, representing four possible directions: up, down, left, and right. Consequently, the urn model must be extended to include a transformation operation $\mathbb{T}$ that acts between four distinct urns, with each site in the two-dimensional lattice hosting four urns. By carefully extending these ideas, the urn model can be generalized to arbitrary higher-dimensional graphs or even complex networks, where each node may be associated with multiple urns corresponding to the available directions or connections. So, this approach offers a versatile and intuitive framework for simulating a broad range of quantum dynamical systems. Our future work will focus on formalizing this methodology and exploring its application to increasingly complex topologies.
The strength of the present model lies in its simplicity and the valuable insights it offers. Successful extensions of this approach could enable the design of DTQW-based algorithms through purely urn-based reasoning and may open new pathways for the efficient quantum simulation of complex classical systems.
Quantum walks have found successful applications across a wide range of domains, including quantum search algorithms [23], community detection in complex networks [24], modeling energy transfer in photosynthesis [25, 26, 27], modeling quantum mechanics in curved space-time [28, 29], to name a few. The split-step quantum walk is particularly notable for its ability to simulate relativistic quantum dynamics [30, 31]. Discrete-time quantum walks have also proven to be powerful tools for simulating topological insulators [32, 33], enabling the study of topologically protected edge states and phase transitions through precise control over coin and shift parameters. Non-Hermitian [34] and open quantum walks extends the standard quantum walk framework to incorporate loss, gain, and environmental interactions, making them powerful tools for modeling quantum active particles [35] and non-equilibrium dynamics in open quantum systems. The present study establishes a connection between the urn model and all these domains where quantum walks are effectively applied.
Properties such as the first-passage problem, stochastic resetting and time of arrival- traditionally studied in the context of classical stochastic processes, can also be explored in quantum walks [36, 37, 38] and other discrete-time quantum processes [39]. Such attempts have been made in the literature using the tools of quantum mechanics. We believe the present work will offer a more accessible framework and open new avenues for future exploration in this direction using the language and tools of the urn model and the corresponding random walks of the marked ball.
In the literature, numerous efforts have been undertaken to simulate quantum behavior using classical frameworks. Broadly, these attempts can be categorized into two approaches: in one, wave functions guide particle trajectories [12]; in the other, wave functions play no dynamical role [13, 14, 15, 16, 17, 18, 40]. Our approach aligns with the latter. The simplicity and accessibility of the framework presented in this study may offer a promising contribution toward foundational insights into the deeper structures underlying quantum theory.
The author thanks Nicolas Gisin, Parongama Sen, and Arghya Das for valuable discussions.
Appendix:
Suppose In the $\mathbb{P}$ process, we set the parameters to match the quantum model: $\eta_{i}=\phi_{i}$ and $\rho_{0}=r_{0}^{2}$. Consequently, the urn occupation number $N_{0}$ is chosen randomly as either
$N_{0}=\lfloor r_{0}^{2}N\rfloor\quad\text{or}\quad N_{0}=\lceil r_{0}^{2}N\rceil$
with equal probability $1/2$. Given that the total number of balls is $N$, the complement is
$N_{1}=N-N_{0}.$
We can always write:
$r_{0}^{2}N=\lfloor r_{0}^{2}N\rfloor+\epsilon=\lceil r_{0}^{2}N\rceil-\epsilon^{\prime},$
where $10\leq\epsilon,\epsilon^{\prime}<$ 1. Moreover, if $r_{0}^{2}N\notin\mathbb{Z}$, then $\epsilon+\epsilon^{\prime}=1$. This implies:
$N_{0}=r_{0}^{2}N+O(1),\quad N_{1}=r_{1}^{2}N+O(1).$
Multiplying gives:
$N_{0}N_{1}=(r_{0}^{2}N+O(1))(r_{1}^{2}N+O(1))=r_{0}^{2}r_{1}^{2}N^{2}+O(N).$
We write $O(f(N))$ to indicate any quantity bounded in absolute value by $Cf(N)$ for some constant $C>0$.
Let us define:
$A:=r_{0}^{2}r_{1}^{2}N^{2},\quad B:=O(N),$
so that:
$\sqrt{N_{0}N_{1}}=\sqrt{A+B}.$
Now factor out the leading-order part:
$\sqrt{A+B}=\sqrt{A}\cdot\sqrt{1+\frac{B}{A}}.$
Since $B=O(N)$ and $A=O(N^{2})$, we have:
$\frac{B}{A}=O\left(\frac{1}{N}\right).$
Using a Taylor expansion:
$\sqrt{1+\frac{B}{A}}=1+\frac{1}{2}\cdot\frac{B}{A}+O\left(\frac{1}{N^{2}}\right),$
which gives:
$\sqrt{A+B}=\sqrt{A}\left(1+O\left(\frac{1}{N}\right)\right)=r_{0}r_{1}N+O(1).$
From the Hadamard-like transformation, we know:
${r_{0}^{\prime}}^{2}=\frac{1}{2}\left(r_{0}^{2}+r_{1}^{2}+2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})\right),$
so multiplying by $N$:
${r_{0}^{\prime}}^{2}N=\frac{1}{2}\left(r_{0}^{2}N+r_{1}^{2}N+2r_{0}r_{1}N\cos(\phi_{0}-\phi_{1})\right).$
Substituting $N_{0}=r_{0}^{2}N+O(1)$, $N_{1}=r_{1}^{2}N+O(1)$, and $\sqrt{N_{0}N_{1}}=r_{0}r_{1}N+O(1)$, we obtain:
${r_{0}^{\prime}}^{2}N=\frac{1}{2}\left(N_{0}+N_{1}+2\sqrt{N_{0}N_{1}}\cos(\phi_{0}-\phi_{1})\right)+O(1).$ (31)
If we define:
$\tilde{N}_{0}:=\frac{1}{2}\left(N_{0}+N_{1}+2\sqrt{N_{0}N_{1}}\cos(\phi_{0}-\phi_{1})\right),$
then:
${r_{0}^{\prime}}^{2}N=\tilde{N}_{0}+O(1).$
Hence, one can equivalently write:
${r_{0}^{\prime}}^{2}N=\lfloor\tilde{N_{0}}\rfloor+O(1)$ (32)
${r_{0}^{\prime}}^{2}N=\lceil\tilde{N_{0}}\rceil+O(1)$ (33)
${r_{0}^{\prime}}^{2}N=\frac{\lfloor\tilde{N_{0}}\rfloor+\lceil\tilde{N_{0}}\rceil}{2}+O(1)$ (34)
Thus, in the asymptotic limit:
$\lim_{N\to\infty}\frac{\lfloor\tilde{N}_{0}\rfloor+\lceil\tilde{N}_{0}\rceil}{2N}={r_{0}^{\prime}}^{2}.$ (35)
Using a similar approach, one can show that in the limit $N\to\infty$, $\eta_{i}^{\prime}(N_{0},N_{1},\phi_{0},\phi_{1})$ of Eq.21 approaches to $\phi_{i}^{\prime}$. Also one can prove using $\tilde{N_{0}}$ and $\eta_{i}^{\prime}$ of Eq. 24 and 26 respectively that the probability of obtaining 0 and 1 in the urn model $\mathbb{M}$ process approaches corresponding probabilities in Quantum version for generalised $SU(2)$ coin operation.